OpenStudy (anonymous):

How do I do quadratic formula on 2x^2 – 7x - 3 = 0

jimthompson5910 (jim_thompson5910):

Plug a = 2, b = -7 and c = -3 into the formula below and simplify \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(-7)\pm\sqrt{(-7)^2-4(2)(-3)}}{2(2)}\] \[\Large x = \frac{7\pm\sqrt{49-(-24)}}{4}\] I'll let you finish